Two straight and infinite long wires separated by a distance “r” carrying current “I” in the same direction will have magnetic field zero then what will be the value of “r”?

Grade:12th pass

1 Answers

Susmita
5 years ago
By any information you provided you cannot find out r.But you can find out the distance where the resultant field will be zero.
Let it be at a distance x from the first wire.
So from the 2nd wire it will be at a distance (r-x).
Resultant magnetic field
B=\frac{\mu_{0}}{2\pi}[\frac{1}{x}-\frac{1}{r-x}]
B=0
Or,x=r-x
Or,2x=r
Or,x=r/2
So halfway between the wire the resultant field is zero.
Please approve if helped.